OpenStudy (anonymous):

Find an equation of g(x) from the table. I know its divided by 2 but I don't know how to write an equation for it. I know I have to use the f(x)=32 but I don't know what to write after that. f(x) 32 16 8 4 2

OpenStudy (nayef):

Which table?

OpenStudy (anonymous):

I am given a table of f(x), g(x), h(x), k(x) but I only need to write an equation for f(x) which is the numbers listed above for f(x) sorry I put g(x) on accident. I know the pattern is each number is divided by 2 so I need to write an equation for that. I just don't know what to put for division.

OpenStudy (anonymous):

unless I put multiplied by 1/2 such as f(x)=32(1/2)^x

OpenStudy (anonymous):

sorry \[f(x) = 2^{x}\]

OpenStudy (anonymous):

2^1= 2 2^2 = 4 2^3 = 8 2^4 = 16 2^5 = 32

OpenStudy (anonymous):

hmm.. so i don't include 32?

OpenStudy (anonymous):

why do you need to include 32? question requires to generate an equation that satisfy the table.

OpenStudy (anonymous):

oh haha Ive never used or generated an equation like before from a table usually i use the number but now I know I can do that too so thanks!

OpenStudy (anonymous):

My pleasure :)

OpenStudy (phi):

to write the equation, we need to know (x, f(x) ) pairs. Your list of numbers in the question is not enough. Are those f(x)? if so, what is x? if the list were x f(x) 0 32 1 16 2 8 3 4 4 2 we could write an equation. We first find the differences between f(x) values: 16, 8, 4 ... These are not constant, so the function is not linear. try finding the ratio between f(x) values: 16/32= 1/2, 8/16= 1/2, etc... We get a ratio of 1/2. So the function is "geometric". Expect a function of the form f(x)= A r^x with r= 1/2 when x is 0 we want f(x)= 32. put those numbers into \[ f(x)= A r^x \\ 32 = A \left( \frac{1}{2}\right)^0 \\ 32=A\] and we get \[ f(x) = 32 \left( \frac{1}{2}\right)^x\] Of course, all of this depends on what the x values are in your table. If we started with x=1, we would have to tweak this answer. If we want to get cute, we can notice 32= 2^5 and (1/2)^x is 2^(-x) , so that \[ f(x) = 32 \left( \frac{1}{2}\right)^x = 2^5 \ 2^{-x} \\ f(x)= 2^{5-x} \] as an alternate way to write the function.